Find the mean of the probability distribution of a random variable X which can take



only the values 1, 2, and 3, given that P(1)=10/33, P(2)= 1/3, and P(3)=1/33

1
Expert's answer
2022-03-23T16:01:20-0400

"\mu=1*\frac{10}{33}+2*\frac{1}{3}+3*\frac{12}{33}=\frac{68}{33}\approx2.0606."

P(3) must be 12/33 not 1/33